# Monomial web basis for the SL(N) skein algebra of the twice punctured sphere

### v1

@@ Abstract.
For any non-zero complex number $q$, excluding finitely many roots of unity of small order, a linear basis for the $\mathrm{SL}(n)$ skein algebra of the twice punctured sphere is constructed. In particular, the skein algebra is a commutative polynomial algebra in $n-1$ generators, where each generator is represented by an explicit $\mathrm{SL}(n)$ web, without crossings, on the surface. This includes the case $q=1$, where the skein algebra is identified with the coordinate ring of the $\mathrm{SL}(n)$ character variety of the twice punctured sphere. The proof of both the spanning and linear independence properties of the basis depends on the so-called $\mathrm{SL}(n)$ quantum trace map, due originally to Bonahon–Wong in the case $n=2$. A consequence of the proof is that the polynomial algebra sits as a distinguished subalgebra of the Lê–Sikora $\mathrm{SL}(n)$ stated skein algebra of the annulus. We end by discussing the relationship with Fock–Goncharov duality.
@@
To a closed oriented three-dimensional manifold $\mathfrak{M}$ containing a, possibly empty, framed oriented knot, or link, $K$ colored by a representation of the Lie group $\mathrm{SU}(n)$, and to a natural number $k$ called the quantization parameter or level, Witten [71] and later Reshetikhin–Turaev [63] assigned a numerical quantum invariant $Z$ now called the WRT invariant. This number $Z$ is expressed in terms of a root of unity $q$ of order approximately $n+k$. The WRT invariant satisfies certain local relations, called skein relations, and when $\mathfrak{M}=\mathfrak{S}^{3}$ is the sphere $Z$ essentially recovers the Jones–Reshetikhin–Turaev $\mathrm{SU}(n)$ knot polynomial associated to the colored knot $K$. For a general manifold $\mathfrak{M}$, the computation of $Z$ bootstraps from the sphere case by using the fact that any such $\mathfrak{M}$ can be obtained by integral Dehn surgery on a link, equivalently linked solid tori, in the sphere, and not to be confused with the link $K$.
@@
Reshetikhin–Turaev, also [67], put Witten’s construction on a more solid mathematical footing by defining the WRT invariant in terms of the representation theory of the quantum group $\mathrm{SU}(n)_{q}$. Lickorish [49] formulated a more elementary and combinatorial approach to computing the $n=2$ invariant by purely skein theoretic methods. One fundamental property employed to this end is that the skein algebra $\mathcal{S}_{2}^{q}(\mathfrak{A})$ of the annulus is isomorphic to the polynomial algebra $\mathbb{C}[x]$ in one variable, allowing for geometrically rich families of polynomials, such as Chebyshev polynomials, to be represented and manipulated pictorially. Analogous skein constructions for the $n=3$ and $n=n$ invariants were developed in [56] and [72], respectively, where the former took advantage of the property that the skein algebra $\mathcal{S}_{3}^{q}(\mathfrak{A})$ is isomorphic to the two variable polynomial algebra $\mathbb{C}[x,y]$.
Main Theorem.
@@
See Theorem 2.1 for a more precise statement. Note it is not claimed that the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{A})$ is not a polynomial algebra for the excluded roots of unity $q$, the study of which is left for the future. However, it is shown that for these excluded $q$ at least one of the generators $x_{i}$ in the main theorem is equal to 00.
To be more precise, there are multiple versions of the $\mathrm{SU}(n)$, or $\mathrm{SL}(n)$, skein algebra, including Sikora’s version described in the main theorem, some of which are briefly discussed now. Based on knotted loops, the HOMFLYPT [22, 60] and $n=2$ Kauffman [37] skein algebras $\mathcal{S}(\mathfrak{S})_{\mathrm{HOMFLYPT}}$ and $\mathcal{S}_{2}^{q}(\mathfrak{S})$ for thickened surfaces $\mathfrak{S}\times(0,1)$, and more generally skein modules for three-dimensional manifolds $\mathfrak{M}$, were initially studied by Turaev [68] and Przytycki [58]. The, in general non-commutative, algebra structure is by superposition relative to the surface. Kuperberg [43] introduced the $n=3$ skein algebra $\mathcal{S}_{3}^{q}(\mathfrak{S})$ based on certain trivalent graphs, called $3$-webs, in addition to loops. Sikora [66] generalized Kuperberg’s algebra to the $n\geq 2$, sometimes written $n=n$, skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$ by incorporating $n$-valent webs. The skein algebra $\mathcal{S}^{q}_{n}(\mathfrak{S})_{\mathrm{CKM}}$ of Cautis–Kamnitzer–Morrison [7] is based on colored trivalent graphs. Lê–Sikora’s [45, 46] stated skein algebra $\mathcal{S}^{q}_{n}(\mathfrak{S})_{\mathrm{st}}$ is similar to Sikora’s skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$, but is sensitive to when the surface has nonempty boundary. In particular, the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{A})$ of the closed annulus equals that of the open annulus, equivalently the twice punctured sphere, but the associated stated skein algebras are different.
It is well-known [66] that the Sikora skein algebra is isomorphic to the CKM skein algebra for $n\leq 3$. As the present paper developed, it was proved by Poudel [57] that such an isomorphism also exists for $n>3$ for all $q$ except possibly for roots of unity of small order. Queffelec–Rose [61] proved, by categorical methods, that the CKM skein algebra of the annulus is a polynomial algebra in $n-1$ generators. The proof of the main result of the present paper has the advantage of being more connected to the underlying cluster geometry discussed later in this introduction, and also identifies a distinguished subalgebra of the stated skein algebra of the annulus isomorphic to the polynomial algebra $\mathbb{C}[x_{1},x_{2},\dots,x_{n-1}]$.
@@
On the other hand, as indicated above, the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$ is more closely tied to the geometry and topology of character varieties. More precisely, when $q=1$ then $\mathcal{S}_{n}^{1}(\mathfrak{S})$ is isomorphic to the unreduced coordinate ring of the $\mathrm{SL}(n)$ character variety of $\mathfrak{S}$ [6, 58, 65]. This means that the reduced coordinate ring is the quotient $\mathcal{S}_{n}^{1}(\mathfrak{S})/\sqrt{0}$ by the set of nilpotent elements. The isomorphism sends a loop, or web, to the associated trace function on the character variety. In particular, the proof of the main theorem gives a new proof that the coordinate ring of the $\mathrm{SL}(n)$ character variety of the annulus is reduced.
@@
The skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$ is naturally a quotient of the $q$-evaluated HOMFLYPT skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})_{\mathrm{HOMFLYPT}}$. Said in another way, multiloops span $\mathcal{S}_{n}^{q}(\mathfrak{S})$ but are not linearly independent. For instance, essentially by the main theorem, Turaev’s loop basis for the HOMFLYPT skein algebra of the annulus becomes linearly dependent in $\mathcal{S}_{n}^{q}(\mathfrak{A})$; compare Lemma 3.12 and Equation ($\ast$3.4). From the point of view of character varieties when $q=1$, this dependence reflects that there are non-trivial relations among trace functions of loops, best captured pictorially in terms of webs [65]. Alternatively, when $q$ is not a root of unity of small order, then $\mathcal{S}_{n}^{q}(\mathfrak{S})$ is spanned by planar webs, namely webs without crossings; see Figure 5. The basis elements for $\mathcal{S}_{n}^{q}(\mathfrak{A})$ of the main theorem are planar webs, thus providing a solution to the problem of finding a linearly independent spanning subset of planar webs.
@@
It is a non-trivial problem to find explicit small generating sets of skein algebras [2, 23], and harder still to find presentations. Indeed, the latter has only been accomplished for a handful of surfaces; see [9, 51] for some more recent progress. The main theorem provides a finite presentation for the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{A})$ of the annulus.
@@
Constructing explicit linear bases of skein algebras is another difficult problem. For small values of $n$ bases have long been known [36, 43], namely the basis of simple essential multi-curves for $n=2$ and of non-elliptic webs for $n=3$, but bases have been more elusive for $n>3$ even for the simplest surfaces; see however [21, 25, 61]. More recently, ideas from hyperbolic and cluster geometry have led to a general blueprint for constructing bases for $\mathcal{S}_{n}^{q}(\mathfrak{S})$, at least when the surface $\mathfrak{S}$ is punctured. When $n$ is $2$ or $3$ this blueprint reproduces the previously mentioned bases, and the present paper can be considered as a demonstration of concept in the case $n>3$ when the surface is the annulus, or twice punctured sphere.
@@
In more detail, such a blueprint is motivated by the concept of Fock–Goncharov duality [18, 19, 30] or, rather, Fock–Goncharov–Shen duality [28, 29] in the current setting. Roughly speaking, this says that there is a canonical linear basis of the coordinate ring of the $\mathrm{SL}(n)$ character variety whose elements are naturally indexed by a discrete sub-monoid $\mathcal{C}_{\mathfrak{L}}\subset\mathbb{N}^{N}$, the positive integer tropical points, defined with respect to an ideal triangulation $\mathfrak{L}$ of the punctured surface $\mathfrak{S}$, and where $N$ is the dimension of the character variety. Here, naturality means that this indexing behaves well under changing the triangulation, in particular transforming according to a tropical cluster mutation formula. Such canonical bases have strong positivity properties; see for example Section 4.4.6, which relates to the physical point of view of Gaiotto–Moore–Neitzke [26]. There is also a quantum version of this duality [12, 20]. From the skein theoretic perspective, being guided by the $n=2$ and $n=3$ cases, the most optimistic scenario when $n>3$ is that such bases are realized by planar webs that can be constructed by a local to global process, first constructing the webs locally in triangles, then gluing together these local webs across the triangulation to obtain the global basis web. This is the procedure followed in the proof of the main theorem.
@@
The key to a skein theoretic realization of Fock–Goncharov duality is the quantum trace map [5, 8, 14, 24, 38, 47, 54]. This is an algebra homomorphism relating two quantizations of the character variety, namely the skein algebra and the Chekhov–Fock–Goncharov quantum higher Teichmüller space. More precisely, it is a map $\tau^{q}_{\mathfrak{L}}$ from the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$ to a quantum torus $\mathcal{T}_{n}^{q}(\mathfrak{S},\mathfrak{L})$, or a $q$-deformed Laurent polynomial algebra in $N$ variables, associated to the ideally triangulated punctured surface $\mathfrak{S}$. Manifesting the tropical geometric nature of Fock–Goncharov duality, the highest term of the quantum trace polynomial of a basis web should have exponents given by the positive tropical integer points of that web. This is true when $n$ is $2$ or $3$, and it is shown to be true when $n>3$ for the annulus in the course of proving the main theorem. In particular, this is the fundamental property underlying the linear independence of the basis elements.
@@
As was somewhat unexpected to the authors, the quantum trace map also plays an important role in establishing the spanning property of the annulus basis, as the inadmissible values of $q$ in the main theorem include the, a priori, unknown roots of a family of polynomials that are forced to be roots of unity of small order by the quantum trace analysis; see Section 3.7.
As an application of the proof of the main theorem, the quantum trace map for the annulus is injective, which is known for all surfaces for $n\leq 3$. In particular, the quantum trace map furnishes an algebra embedding of the polynomial algebra $\mathbb{C}[x_{1},x_{2},\dots,x_{n-1}]$ into a, non-commutative, quantum torus $\mathcal{T}_{n}^{q}(\mathfrak{A},\mathfrak{L})$ for every admissible $q$. See Section 5.1. Another closely related application is the injectivity of the splitting map [45, 46] for the stated skein algebra $\mathcal{S}^{q}_{n}(\mathfrak{A})_{\mathrm{st}}$ of the twice punctured sphere for the admissible $q$ of the main theorem. For general punctured surfaces, with empty boundary, this is known for all $q$ when $n\leq 3$ [34, 45], and when $n>3$ for $q=1$ [42, 69] and for certain roots of unity $q$ of small order [70]. See Section 5.2.

### v2

@@ Abstract.
We give a new proof of a slightly modified version of a result of Queffelec–Rose, by constructing a linear basis for the $\mathrm{SL}(n)$ skein algebra of the twice punctured sphere for any non-zero complex number $q$, excluding finitely many roots of unity of small order. In particular, the skein algebra is a commutative polynomial algebra in $n-1$ generators, where each generator is represented by an explicit $\mathrm{SL}(n)$ web, without crossings, on the surface. This includes the case $q=1$, where the skein algebra is identified with the coordinate ring of the $\mathrm{SL}(n)$ character variety of the twice punctured sphere. The proof of both the spanning and linear independence properties of the basis depends on the so-called $\mathrm{SL}(n)$ quantum trace map, due originally to Bonahon–Wong in the case $n=2$. Two consequences of our method are that the quantum trace map and the so-called splitting map embed the polynomial algebra into the Fock–Goncharov quantum higher Teichmüller space and the Lê–Sikora stated skein algebra, respectively, of the annulus. We end by discussing the relationship with Fock–Goncharov duality.
@@
To a closed oriented three-dimensional manifold $\mathfrak{M}$ containing a, possibly empty, framed oriented knot, or link, $K$ colored by a representation of the Lie group $\mathrm{SU}(n)$, and to a natural number $k$ called the quantization parameter or level, Witten [76] and later Reshetikhin–Turaev [66] assigned a numerical quantum invariant $Z$ now called the WRT invariant. This number $Z$ is expressed in terms of a root of unity $q$ of order approximately $n+k$. The WRT invariant satisfies certain local relations, called skein relations, and when $\mathfrak{M}=\mathfrak{S}^{3}$ is the sphere $Z$ essentially recovers the Jones Reshetikhin–Turaev $\mathrm{SU}(n)$ knot polynomial associated to the colored knot $K$. For a general manifold $\mathfrak{M}$, the computation of $Z$ bootstraps from the sphere case by using the fact that any such $\mathfrak{M}$ can be obtained by integral Dehn surgery on a link, equivalently linked solid tori, in the sphere, and not to be confused with the link $K$.
@@
Reshetikhin–Turaev, also [72], put Witten’s construction on a more solid mathematical footing by defining the WRT invariant in terms of the representation theory of the quantum group $\mathrm{SU}(n)_{q}$. Lickorish [52] formulated a more elementary and combinatorial approach to computing the $n=2$ invariant by purely skein theoretic methods. One fundamental property employed to this end is that the skein algebra $\mathcal{S}_{2}^{q}(\mathfrak{A})$ of the annulus is isomorphic to the polynomial algebra $\mathbb{C}[x]$ in one variable, allowing for geometrically rich families of polynomials, such as Chebyshev polynomials, to be represented and manipulated pictorially. Analogous skein constructions for the $n=3$ and $n\geq 2$, sometimes written $n=n$, invariants were developed in [59] and [78], respectively, where the former took advantage of the property that the skein algebra $\mathcal{S}_{3}^{q}(\mathfrak{A})$ is isomorphic to the two variable polynomial algebra $\mathbb{C}[x,y]$.
The primary contribution of the present paper is a new proof of a slightly modified version of the following result.
Theorem (Queffelec–Rose [64]).
The $\mathrm{SU}(n)$ skein algebra of the annulus is isomorphic to a polynomial algebra $\mathbb{C}[x_{1},x_{2},\dots,x_{n-1}]$ in $n-1$ generators.
More precisely, we prove the following result.
Theorem 1.
@@
See Theorem 2.1 for the most precise statement.
To be more clear, there are multiple versions of the $\mathrm{SU}(n)$, or $\mathrm{SL}(n)$, skein algebra, including Sikora’s version described in Theorem 1, some of which are briefly discussed now. Based on knotted loops, the HOMFLYPT [23, 63] and $n=2$ Kauffman [38] skein algebras $\mathcal{S}(\mathfrak{S})_{\mathrm{HOMFLYPT}}$ and $\mathcal{S}_{2}^{q}(\mathfrak{S})$ for thickened surfaces $\mathfrak{S}\times(0,1)$, and more generally skein modules for three-dimensional manifolds $\mathfrak{M}$, were initially studied by Turaev [73] and Przytycki [61]. The, in general non-commutative, algebra structure is by superposition relative to the surface. Kuperberg [45] introduced the $n=3$ skein algebra $\mathcal{S}_{3}^{q}(\mathfrak{S})$ based on certain trivalent graphs, called $3$-webs, in addition to loops. Sikora [69] generalized Kuperberg’s algebra to the $n=n$ skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$ by incorporating $n$-valent webs. The skein algebra $\mathcal{S}^{q}_{n}(\mathfrak{S})_{\mathrm{CKM}}$ of Cautis–Kamnitzer–Morrison [7] is based on colored trivalent graphs, and is the algebra described in the above theorem by Queffelec–Rose. Lê–Sikora’s [47, 49] stated skein algebra $\mathcal{S}^{q}_{n}(\mathfrak{S})_{\mathrm{st}}$ is similar to Sikora’s skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$, but is sensitive to when the surface has nonempty boundary. In particular, the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{A})$ of the closed annulus equals that of the open annulus, equivalently the twice punctured sphere, but the associated stated skein algebras are different.
Note it is not claimed in Theorem 1 that the Sikora skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{A})$ is not a polynomial algebra for the excluded roots of unity $q$, and it is indeed true that $\mathcal{S}_{n}^{q}(\mathfrak{A})$ is a polynomial algebra for $n=2,3$ for all $q$. However, it is shown that for these excluded $q$ at least one of the generators $x_{i}$ in Theorem 1 is equal to 00 in the skein algebra.
It is well-known [69] that the Sikora skein algebra is isomorphic to the CKM skein algebra for $n\leq 3$. As the present paper developed, it was proved by Poudel [60] that such an isomorphism also exists for $n>3$ for all $q$ except possibly for roots of unity of small order. Queffelec–Rose’s proof of the above theorem was by categorical and representation theoretic methods; see also [65]. Moreover, their result is valid for any $q$, or more generally for the skein algebra viewed as a $\mathbb{Z}[q,q^{-1}]$-module. Our proof of Theorem 1 has the advantage of being more connected to the underlying cluster geometry discussed later in this introduction.
@@
On the other hand, as indicated above, the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$ is more closely tied to the geometry and topology of character varieties. More precisely, when $q=1$ then $\mathcal{S}_{n}^{1}(\mathfrak{S})$ is isomorphic to the unreduced coordinate ring of the $\mathrm{SL}(n)$ character variety of $\mathfrak{S}$ [6, 61, 68]. This means that the reduced coordinate ring is the quotient $\mathcal{S}_{n}^{1}(\mathfrak{S})/\sqrt{0}$ by the set of nilpotent elements. The isomorphism sends a loop, or web, to the associated trace function on the character variety. In particular, the proof of Theorem 1 gives another proof that the coordinate ring of the $\mathrm{SL}(n)$ character variety of the annulus is reduced. See also [18, 70].
@@
The skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$ is naturally a quotient of the $q$-evaluated HOMFLYPT skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})_{\mathrm{HOMFLYPT}}$. Said in another way, multiloops span $\mathcal{S}_{n}^{q}(\mathfrak{S})$ but are not linearly independent. For instance, essentially by Theorem 1, Turaev’s loop basis for the HOMFLYPT skein algebra of the annulus becomes linearly dependent in $\mathcal{S}_{n}^{q}(\mathfrak{A})$; compare Lemma 3.12 and Equation ($\ast$3.4). From the point of view of character varieties when $q=1$, this dependence reflects that there are non-trivial relations among trace functions of loops, best captured pictorially in terms of webs [68]. Alternatively, when $q$ is not a root of unity of small order, then $\mathcal{S}_{n}^{q}(\mathfrak{S})$ is spanned by planar webs, namely webs without crossings; see Figure 5. The basis elements for $\mathcal{S}_{n}^{q}(\mathfrak{A})$ of Theorem 1 are planar webs, thus providing a solution to the problem of finding a linearly independent spanning subset of planar webs.
@@
It is a non-trivial problem to find explicit small generating sets of skein algebras [2, 24], and harder still to find presentations. Indeed, the latter has only been accomplished for a handful of surfaces; see [9, 54] for some more recent progress. Theorem 1 provides a finite presentation for the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{A})$ of the annulus.
@@
Constructing explicit linear bases of skein algebras is another difficult problem. For small values of $n$ bases have long been known [37, 45], namely the basis of simple essential multi-curves for $n=2$ and of non-elliptic webs for $n=3$, but bases have been more elusive for $n>3$ even for the simplest surfaces; see however [22, 26, 64]. More recently, ideas from hyperbolic and cluster geometry have led to a general blueprint for constructing bases for $\mathcal{S}_{n}^{q}(\mathfrak{S})$, at least when the surface $\mathfrak{S}$ is punctured. When $n$ is $2$ or $3$ this blueprint reproduces the previously mentioned bases, and the present paper can be considered as a demonstration of concept in the case $n=n$ when the surface is the annulus.
@@
In more detail, such a blueprint is motivated by the concept from mirror symmetry of Fock–Goncharov duality [19, 20, 31] or, rather, Fock–Goncharov–Shen duality [29, 30] in the current setting. Roughly speaking, this says that there is a canonical linear basis of the coordinate ring of the $\mathrm{SL}(n)$ character variety of $\mathfrak{S}$ whose elements are naturally indexed by a discrete sub-monoid $\mathcal{C}_{\mathfrak{L}}\subset\mathbb{N}^{N}$, the positive integer tropical points, defined with respect to an ideal triangulation $\mathfrak{L}$ of the punctured surface $\mathfrak{S}$, and where $N$ is the dimension of the character variety. Here, naturality means that this indexing behaves well under changing the triangulation, in particular transforming according to a tropical cluster mutation formula. Such canonical bases have strong positivity properties; see for example Section 4.4.6, which relates to the physical point of view of Gaiotto–Moore–Neitzke [27]. There is also a quantum version of this duality [12, 21]. From the skein theoretic perspective, being guided by the $n=2$ and $n=3$ cases, the most optimistic scenario when $n=n$ is that such bases are realized by planar webs that can be constructed by a local to global process, first constructing the webs locally in triangles, then gluing together these local webs across the triangulation to obtain the global basis web. This is the procedure followed in the proof of Theorem 1.
@@
The key to a skein theoretic realization of Fock–Goncharov duality is the $\mathrm{SL}(n)$ quantum trace map [5, 8, 14, 25, 39, 50, 57]. This is an algebra homomorphism relating two quantizations of the character variety, namely the skein algebra and the Chekhov–Fock–Goncharov quantum higher Teichmüller space. More precisely, it is a map $\tau^{q}_{\mathfrak{L}}$ from the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$ to a quantum torus $\mathcal{T}_{n}^{q}(\mathfrak{S},\mathfrak{L})$, or a $q$-deformed Laurent polynomial algebra in $N$ variables, associated to the ideally triangulated punctured surface $\mathfrak{S}$. Manifesting the tropical geometric nature of Fock–Goncharov duality, the highest term of the quantum trace polynomial of a basis web should have exponents given by the positive tropical integer points of that web. This is true when $n$ is $2$ or $3$, and it is shown to be true when $n=n$ for the annulus in the course of proving Theorem 1. In particular, this is the fundamental property underlying the linear independence of the basis elements.
@@
As was somewhat unexpected to the authors, the quantum trace map also plays an important role in our proof in establishing the spanning property of the annulus basis. This is because the inadmissible values of $q$ in Theorem 1 include the, a priori, unknown roots of a family of polynomials. These roots are forced to be roots of unity of small order by the quantum trace analysis; see Section 3.7.
As a consequence of the proof of Theorem 1, we establish the following parallel applications relating quantum algebras associated to the annulus $\mathfrak{A}$.
Theorem 2.
The quantum trace map for the skein algebra of the annulus is injective for every admissible value of $q$ as in Theorem 1. In particular, the quantum trace map furnishes an algebra embedding of the polynomial algebra $\mathbb{C}[x_{1},x_{2},\dots,x_{n-1}]$ into a quantum torus $\mathcal{T}_{n}^{q}(\mathfrak{A},\mathfrak{L})$ depending on the choice of an ideal triangulation $\mathfrak{L}$ of the annulus.
See Corollary 5.1 in Section 5.1. Note that the injectivity of the quantum trace map is known for general punctured surfaces $\mathfrak{S}$ when $n\leq 3$ for all $q$.
Theorem 3.
The splitting map [47, 49] for the skein algebra of the annulus is injective for every admissible value of $q$ as in Theorem 1. In particular, the splitting map furnishes a canonical algebra embedding of the polynomial algebra $\mathbb{C}[x_{1},x_{2},\dots,x_{n-1}]$ into the stated skein algebra $\mathcal{S}^{q}_{n}(\mathfrak{A})_{\mathrm{st}}$.
See Corollary 5.4 in Section 5.2. Note that the injectivity of the splitting map is known for general punctured surfaces $\mathfrak{S}$ when $n\leq 3$ for all $q$ [35, 47] and when $n=n$ for $q=1$ [44, 74] and for certain roots of unity $q$ of small order [75].